### v1

# Medium Access for Push-Pull Data Transmission in 6G Wireless Systems

Abstract
Medium access in 5G systems was tailored to accommodate diverse traffic classes through network resource slicing. 6G wireless systems are expected to be significantly reliant on Artificial Intelligence (AI), leading to data-driven and goal-oriented communication. This leads to augmentation of the design space for Medium Access Control (MAC) protocols, which is the focus of this article. We introduce a taxonomy based on push-based and pull-based communication, which is useful to categorize both the legacy and the AI-driven access schemes. We provide MAC protocol design guidelines for pull- and push-based communication in terms of goal-oriented criteria, such as timing and data relevance. We articulate a framework for co-existence between pull and push-based communications in 6G systems, combining their advantages. We highlight the design principles and main tradeoffs, as well as the architectural considerations for integrating these designs in Open-Radio Access Network (O-RAN) and 6G systems.
Index Terms:
IoT connectivity, pull-based communication, push-based communication, medium access control, timing, data relevance
I Introduction
One of the major promises of the sixth generation of mobile networks (6G) is a seamless integration of technology, data, and intelligence in real time [1]. This is a key enabler for new applications such as eXtended Reality (XR) and digital twins (DTs), which require a two-way connection between the digital and physical worlds: decisions in the digital domain have real, tangible effects, and real events need to be integrated in the digital model of the environment. Naturally, this level of real-time integration may strain even the capacity of future 6G networks, particularly when there is a massive number of connected devices [2] and the computational requirements of Artificial Intelligence (AI) models are considered.
The novel services envisioned with 6G systems expect flexible communication protocols designed to accommodate different applications and requirements. The network slicing approaches proposed for 5G can create multiple virtual networks tailored to specific service needs on a shared physical infrastructure to meet the requirements of different traffic classes. However, due to the dynamic variability of end-to-end goals over time as well as the use of AI by the communicating parties, 6G systems will require agile and robust Medium Access Control (MAC) protocols to accommodate diverse application-driven requirements, not catered efficiently by the existing network slicing strategies [1]. The communication protocols for 6G systems should be designed considering the growing intelligence in the communicating nodes and the coexistence of diverse application requirements and communication modalities.
The resulting high reliability and flexibility requirements pose two key challenges: first, where to process the information under strict latency-reliability requirements, and secondly, when and how to collect the data for processing. The former aspect has been extensively discussed in the domain of edge computing and task offloading. The latter aligns with the goal-oriented communication paradigm [3], which optimizes communication decisions based on the Value of Information (VoI) of updates: when transmitting all available data is impossible due to timing or energy constraints, goal-oriented schemes privilege information that is both novel and relevant for the receiver, i.e., is both not directly predictable from previously available information and useful for the task at hand [2].
The increased role of AI gave rise to the notions of semantic and goal-oriented communication, focusing on the extraction of relevant information [3]. Goal-oriented optimization of MAC mechanisms entails coordination of multiple agents with a partial view of the system. For example, uncoordinated decisions made by Internet of Things (IoT) devices for uplink transmission are myopic and may deteriorate performance due to the risk of collisions. The majority of existing goal-oriented schemes operate in a pull-based fashion [4]: the Base Station (BS) is directly connected to the DT of the system and is thus able to estimate which information is expected to be more relevant, proactively requesting it from the relevant devices. However, pull-based systems can only rely on statistical information, while the sensors themselves have access to the actual observed information and may have a more precise idea of its value. Ideally, decentralized push-based access may lead to more informative updates and a better use of wireless resources [5]. However, this brings challenges of coordination, as individual sensors have a limited view of the environment, and random access schemes have well-known congestion issues.
Figure 1: A toy example of an IoT scenario where a BS collects measurements from IoT devices to construct a digital twin (DT) of the system. The coexistence of push- and pull-based communication enables the BS to coordinate requests for the most informative data to update the DT model while allowing IoT devices to immediately report detected anomalies.
In fact, the two approaches have complementary benefits: push-based communication follows an event-driven approach, excelling in scenarios that require immediate reporting and response, such as anomaly detection. Pull-based communication is more effective in scenarios where the edge or cloud runs a comprehensive AI model that can impose global orchestration of the data collection and selectively retrieve data, thereby reducing irrelevant data traffic. Moreover, it offers flexibility in adopting communication strategies to handle requests for different tasks by, for example, deciding the timing of data collection to achieve a specific goal [6]. This is evident in the scenario shown in Fig. 1: the uncertainty of the DT model over the robot (node 5) is high, so the BS can initiate a pull request to get that information. In this case, the global view of the system allows the BS to reduce communication by, e.g., exploiting the information from the camera (node 3) to avoid contacting the other robot (node 6) as well. On the other hand, one of the sensors (node 2) is registering an anomaly, which is not reflected in the DT. This event would be detected very late under a purely pull-based scheme, and requires the node to be able to transmit at will.
The dichotomy between classical and goal-oriented schemes then needs to be complemented by a second axis, namely, push- versus pull-based schemes. This distinction is already present in classical schemes, as scheduled mechanisms are pull-based, while random access is push-based. Fig. 2 shows a graphical representation of these aspects: while pull-based schemes rely on centralized knowledge of packet timing, which push-based schemes lack, the distinction between classical and AI-driven semantic schemes [7] hinges on the availability of data, i.e., knowledge about the meaning and value of data that each sensor has, and the awareness of timing for data availability. This adds a novel semantic dimension in the design space of access schemes, as interactions between the BS and the devices are strategic. While classical MAC schemes operate independently from this aspect, it is at the center of goal-oriented communication schemes, which target the VoI of potential updates, privileging more informative updates for the DT of the environment.
The design of protocols and MAC mechanisms for the coexistence of push- and pull-based traffic supporting goal-oriented Key Performance Indicators (KPIs), such as estimation error or control performance, in addition to traditional timing metrics, is a key enabler for real-time applications such as DTs and XR [8]; addressing this gap is imperative to the design of efficient 6G systems. As knowledge about the environment is itself distributed, centralizing decision-making at the BS or distributing it to the sensors leads to highly different information structures and signaling requirements.
This article introduces the design criteria for MAC protocols to accommodate push- and pull-based communications, offering guidelines for organizing intelligent operation in 6G wireless access networks. To that end, we design a time-division mechanism and provide numerical studies in selected use cases to demonstrate the validity of the developed analytical models and their potential for application-driven 6G network operation.
II Goal-Oriented Communication: A MAC View
Figure 2: Design space of push and pull communication paradigms depending on the data availability and timing of arrival.
In a goal-oriented scenario such as the one depicted in Fig. 1, each sensor may have precise knowledge of specific physical parameters based on its own measurements, but this knowledge is limited to the status of a local part of the environment: the sensor has no way to obtain a global view of the environment, other than through a transmission from the BS. Aside from the signaling issues required to maintain and synchronize even a partial copy of the environment DT at each sensor, there are also significant computational limitations: sensors are often simple battery-powered devices with low-power computing units, which cannot run AI models in real time. Pull-based communication seems to be the best possible solution to this issue [6]; since the BS is the only node with a global view of the system, centralizing decisions is an attractive option. Centralized scheduling also mitigates the collision problem, as it allows for orthogonal division of channel resources to avoid interference.
On the other hand, having a global view and easily coordinating access comes at a cost: since it is based on a statistical model, it relies on the expected behavior of the system, and it might not reflect physical reality. Unexpected events, such as anomalies or component failures, might not be accurately modeled, or assigned a very low probability. Only push-based communication, in which individual sensors decide when to transmit based on the actual measurements they make [5], is able to overcome this issue and quickly make the BS aware of the anomaly. However, push-based schemes suffer from coordination problems, negating the advantages of pull-based scheduling.
More complex sampling schemes, such as Semantic Data Sourcing (SEMDAS) [9], allow for content-based or semantic queries, which direct sensors to transmit if their measurements are within the (possibly complex) parameters of the query. This hybrid solution maintains and even enhances the advantages of having a global view of the system, but loses the coordination advantages of purely pull-based solutions, as the number and identities of the selected nodes might be unknown even to the BS itself. However, the fundamental trade-off remains, and the relevant literature is still considering push- and pull-based schemes as mutually exclusive. The coexistence between push- and pull-based access in the same system, enabling both global coordination for normal operation and quick anomaly reporting, is a new challenge that opens new fundamental trade-offs and design problems.
Figure 3: A time-diagram for the considered push-pull frame structures.
Deciding whether to allocate specific resources to push-based operation, and how to balance between pull- and push-based transmission, represents a complex problem that depends on the application requirements. Static, predefined allocation schemes cannot adapt to the conditions of the environment, while dynamic schemes require additional signaling to inform nodes of the resource allocation. Moreover, coordinating push-based transmissions becomes even more complex if nodes are also regularly transmitting data in a pull-based fashion, changing the push-based decision-making due to recent or upcoming pull-based transmission opportunities.
III Protocol design
Combining the push- and pull-based paradigms requires careful protocol designs, accounting for application-level requirements and KPIs. In this section, we highlight the design principles for communication protocols managing the coexistence of push- and pull-based data transmission schemes, along with key trade-offs for two specific goal-oriented scenarios: centralized decision-making and decentralized intelligence.
III-A Centralized decision-making
In the centralized decision-making scenario, the BS controls medium access for both push- and pull-based traffic considering the application-level requirement and expected traffic for each class [10]. Our recent work outlines two types of MAC layer designs, as illustrated in Fig. 3. One is the contention-free pull and contention-push (CFC-pull/push), which consists of two parts: scheduled contention-free slots for pull-based communication and random access contention-based slots for push-based communication [10]. The other is Reserved pull-contention and shared pull-push contention (RCSC-pull/push), which also consists of two parts: reserved random access periods for pull-based traffic and shared random access periods for both traffic classes [11]. In both structures, the total available time slots per frame are organized into corresponding sub-frames by adjusting a control parameter $\alpha\in[0,1]$, representing the fraction of resources reserved for pull-based communication.
CFC-pull/push is a frame structure similar to the contention-free and contention-based mechanism in WiFi technology. Each frame’s organization is notified to the devices through control signaling [10]: firstly, the BS decides the device scheduling for the pull-based communication phase, taking into account the available pull-slots, and notifies devices by broadcasting the scheduling information. This unicast scheduling enables the BS to retrieve sensor readings from specific devices at specific times without packet collisions. Conversely, during the push-contention phase, each device transmits an update – if it deems it is important enough – by contending for the channel with other push-based traffic. The major challenge is to design the resource allocation scheme that maximizes the goal-oriented utility for both phases [10]. Fig. 4 shows the average incoming traffic that can be served by the system with a target success probability as a function of $\alpha$. Every point in the plot represents a different number of slots reserved for responding to pull requests, $\lfloor\alpha S\rfloor$. As expected, increasing this value leads to a decrease in the acceptable push traffic, while accommodating a higher pull traffic, and vice versa. Naturally, reducing the target success probability allows the BS to manage a higher overall load.
[…]
Figure 4: CFC-pull/push achievable incoming traffic satisfying a target transmission success probability $p_{\mathrm{th}}$ as a function of $\alpha$ [10].
Protocol designs based on RCSC-pull/push are crucial when data collection by the BS is based on the specific observations of each device [11]. The RCSC-pull/push can be divided into two portions: pull-reserved contention slots and pull-push shared contention slots. At the beginning of the frame, the BS sends a control signal to push-enabled devices to notify them of the frame structure. The BS then broadcasts a semantic query, which specifies the features of the data that the BS considers relevant [11], e.g., by specifying a range over the value of sensor observations, to all pull-enabled devices. Pull-enabled devices receiving this query check whether their observations satisfy the condition specified by the BS and transmit based on a random access scheme. On the other hand, devices operating in a push-based fashion start transmitting data at the beginning of the shared slots if they observe informative data, even if it is outside the query specifications. Within the shared slots, push-based transmissions need to contend for the channel with the pull-based responses that have not yet been successful by the end of the pull reserved slots. This results in both intra-class and inter-class collisions [11]. Optimizing this frame structure entails trade-offs between different task-specific communication requirements. Fig. 5 illustrates the impact of varying $\alpha$ on the content-driven communication scenario. Performance is evaluated in terms of the pull-based retrieval accuracy and push-based transmission success probability. The former represents the probability that the BS successfully receives all requested content, while the latter is the transmission success probability for pushed data. Fig. 5 shows that increasing $\alpha$ improves pull-based retrieval accuracy, since most sensors are able to successfully transmit in the reserved slots. On the other hand, higher values of $\alpha$ reduce the push-based success probability due to fewer available transmission opportunities and more intra-class collisions.
[…]
Figure 5: RCSC-pull/push performance in terms of pull retrieval accuracy (solid lines) and push success probability (dashed lines) as a function of $\alpha$, and for different number $S$ of slots in a frame.
III-B Decentralized intelligence
In the decentralized intelligence scenario, devices make local observations and estimate their expected value to decide when to transmit. As individual devices have limited knowledge about others’ observations, transmissions are subject to collisions, and thus to the loss of timely and accurate status updates. Devices then need to decide whether and when to attempt a transmission so as to satisfy the KPIs for the goal-oriented tasks, such as VoI [5]. As even the BS has partial information on the value of data at the nodes, this leads to a different choice of MAC frame design and access protocols, as the frame structure needs to be adapted significantly to match the requirements set by the emerging applications and use cases.
For instance, consider a distributed learning setting for collective intelligence. The BS would use private data available at the IoT devices, which have limited computational and communication capabilities, to train high-quality learning models. To do so, the BS can reuse the CFC-pull/push frame structure and the corresponding communication protocol to solicit application-relevant data. This means devising efficient scheduling policies to accelerate high-quality training, where contention-free pull-reserved slots are allocated to the devices with the most useful information in each frame. A wrong attribution of the value of data from specific devices during scheduling, due to the partial knowledge of the BS, can result in a weaker model with poor generalization performance. To overcome this issue, devices that are not accommodated in the pull slots need transmission opportunities to share a better representation of their data. This can be done by arranging available communication resources for pull- and push-based slots, such that a high-quality learning model can be trained in a communication-efficient manner, as shown in Fig. 6. A reference protocol design for such a case is defined in [12].
Furthermore, devices need to balance the urgency of updates that might have to wait for a long time before the next pull-based slot with the relative scarcity of push-based slots and the consequent risk of collisions. Goal-oriented operation might require reasoning about other devices’ beliefs, exploiting publicly available information to come to a consensus without further signaling. Naturally, aside from the computational load that this puts on the nodes, it also requires a careful design of the beacon and acknowledgment messages.
5102030400.880.880.880.880.90.90.90.90.920.920.920.920.940.940.940.94Number of push IoT devices per frameTest AccuracyRandomStrategic push-pullOnly-pull
Figure 6: Learning performance under a fixed number of slots per frame; strategic access push-pull improves accuracy over Only-pull, and outperforms Random scheduling significantly.
IV Integration Challenges and Opportunities
In this section, we identify and discuss key challenges in protocol designs to enable the integration of push-pull data transmission in 6G systems.
IV-A Intelligent MAC protocols
The goal-oriented aspect of communication outlines the value of the data collected at the receiver, instead of quantifying it locally at the device collecting the data, following task-specific requirements. Therefore, the underlying MAC mechanism requires adaptability and flexibility to allow intelligent coordination between IoT devices and the BS. As individual devices are oblivious to the global view, they can only act according to their local knowledge of the value of available data upon transmission when employing conventional push-based schemes. Alternatively, devices may be able to perform a strategic push by inferring the potential value of data transmission through past feedback from the BS, which implicitly reflects the knowledge on the state of the system. This can alleviate the complexity of distributed coordination and intelligence when optimizing MAC. Similarly, the BS can exploit its knowledge of prior interactions and employ a more conservative mechanism in the following frames, performing a strategic pull.
However, the risk of missing out unforeseen critical updates, such as alarms, always exists in the absence of available push communication resources. In practice, the previous considerations require the BS to be able to dynamically modify the MAC structure to achieve different trade-offs, switching between CFC-pull/push and RCSC-pull/push frames and adjusting the number of pull, push, or shared slots (see Fig. 3). This also translates to the need for specific and reliable control signals broadcasting the information regarding the MAC design to all devices in a timely manner. In the extreme case, the MAC structure could be modified on a frame-by-frame basis, but trade-offs between frame flexibility and the overhead caused by the additional control information should be investigated.
IV-B Energy-Efficient Connectivity
Energy-efficient protocol designs for centralized decision-making and decentralized intelligence are crucial to realize energy-efficient connectivity in 6G networks. Technological enablers such as Wake-up Radio (WuR) can be leveraged to design sustainable communication systems and protocols that support intelligent tasks at the network edge [13]. Our recent work [10, 11] outlines communication framework designs incorporating WuR to manage push-pull coexistence, allowing IoT devices to keep their primary transceivers off until they detect a pull request. Specifically, unicast wake-up (UCWu) [14] is introduced for CFC-pull/push to retrieve data from a specific device in a grant-based manner. Conversely, Content-based Wake-up (CoWu) is introduced to enable content-based pull with WuRs in the RCSC-pull/push setting [11]. In CoWu, the information about the desired content is embedded into the Wake-up Signal (WuS), allowing only the nodes with relevant data to activate their primary transceiver.
While employing WuR can improve the overall energy efficiency, it also adds design challenges. Highly informative feedback from the BS, for example, aiming to improve future transmission policy through strategic push, cannot be realistically performed through WuSs due to their low data rate [14]. Furthermore, a dynamic adoption of WuSs within the flexible MAC protocols is needed to tailor the correct push-pull paradigm per the design space, accordingly to the current context of data collection and the goal-oriented scenario (see Fig. 2). Such dynamic adaptation should consider the pros and cons of grant-based and grant-free access in terms of system-level performance, including energy efficiency.
Compute operations needed for AI training and inference add a significant energy load to the IoT networks. In a collaborative learning scenario, the communication overhead for distributed coordination between devices to collect/exchange model updates is significant. With the increasing importance of AI technology in 6G networks, communication protocols should also take into account the energy cost of employing AI models and their interaction with the communication system for goal-oriented tasks. Low-power AI models, such as TinyML and Spiking Neural Networks (SNNs), can serve as technology enablers for understanding data semantics locally and transmitting valuable model updates that improve the learning performance [12]: the BS can use the estimates of data availability, timing, and relevance for scheduling devices’ updates during training, while allowing devices to also share learning models based on their local knowledge.
Specifically, while a BS could retrieve only high-quality data that improves its model accuracy through strategic pull, the devices can strategically push relevant model parameters and data, or even anomalous samples detected locally by the installed lightweight AI models, avoiding unnecessary data transmissions. However, this also indicates the need for further investigation of the coexistence between push- and pull-based access and the design of efficient MAC structures for goal-oriented tasks, such as learning and inference. While countermeasures for handling the excess cost of communication and processing before deployment at the devices, e.g., by exploiting well-known compression techniques such as knowledge distillation, can be enacted, these models should be strategically updated taking into account the available energy, traffic load, and the task-specific performance requirements. Therefore, the underlying MAC mechanisms need to consider AI models and data distribution, as well as the cost of their deployment and exploitation on-device, to fully support energy-efficient connectivity.
IV-C 6G integration
A seamless integration of push-pull data transmissions in future 6G networks necessitates both a higher-layer application able to translate the current goal-oriented requirements and the status of the network into a viable allocation of network resources, and advanced Radio Resource Control (RRC) mechanisms capable of dynamic, real-time radio resource allocation based on the chosen MAC structure. The former can be obtained through the development of Open-Radio Access Network (O-RAN)-compliant xApps/rApps [15], while the latter requires flexible management of the physical (PHY) channels.
O-RAN-Compliant xApps/rApps
xApps operating within the Near-Real-Time RAN Intelligent Controller (Near-RT RIC) can serve as proxies between the entity requesting the data, e.g., the DT, and the next-generation Node B (gNB). On one side, xApps can interpret high-level requirements arriving through the Y1 interface – such as data content preferences and deadlines – and combine this information with network status feedback provided by the gNB via the E2 interface [15]. This enables the conversion of goal-oriented requirements into resource configurations, which are sent back to the gNB for implementation through dynamic RRC. Moreover, xApps are responsible for processing data acquired from the network before transmitting it to the end destination via the Y1 interface. Finally, in scenarios where multiple agents request data with potentially conflicting objectives, an xApp must aggregate and prioritize these requirements to ensure fair and efficient management of network resources. Similar operations can be implemented in rApps operating within the Non-Real-Time RIC (Non-RT RIS) [15], where near-real-time control loops are not needed.
Dynamic RRC
To effectively implement the resource configurations determined by xApps, the gNB must employ RRC to dynamically manage radio resource allocation for each PHY channel, organizing their placement within the MAC frame. In 5G, contention-based communication utilizes the Physical Random Access Channel (PRACH), primarily for initial access. This channel can support contention-based transmission in push-pull MAC designs, i.e., for the push portion of the CFC-pull/push frame and throughout the RCSC-pull/push frame. Conversely, the Physical Uplink Shared Channel (PUSCH) handles scheduled uplink data transmissions, enabling uni-cast pull-based communications within the CFC-pull/push. Moreover, the Physical Downlink Control Channel (PDCCH) can convey control signaling regarding dynamic RRC information, including MAC design decisions and scheduling directives. A flexible allocation of these channels across available PHY resources is crucial to accommodate diverse push-pull data transmission paradigms. Moreover, integrating energy-efficient technologies such as WuR introduces additional complexity in control signaling interactions. For example, to enable the CFC-pull/push frame structure with UCWu, a WuS containing the device ID could be sent via the PDCCH– or a newly defined physical WuS channel – before each pull slot, which in turn utilizes the PUSCH. Then, the PRACH would be employed to implement push slots, potentially occurring with a periodicity lower than $10$ ms – the minimum allowed by the 5G standard – depending on the frame duration [10]. Therefore, exploring more flexible RRC approaches is essential for integrating push-pull coexistence into future 6G networks.
V Conclusion
This work provides MAC protocol design guidelines based on novel pull- and push-based communication schemes for goal-oriented data transmission in 6G systems. We articulate the need to consider timing and data availability as fundamental measures for assessing data relevance and enabling intelligent network operations, such as adaptive and flexible frame structure to support application-level requirements and KPIs. Using examples of goal-oriented scenarios, we present key trade-offs in the coexistence of pull- and push-based communication and postulate a design space for access schemes that includes an additional semantic dimension. We conclude by outlining potential integration challenges and opportunities with the presented push-pull data transmission schemes, particularly when data-driven intelligence and requirements for goal-oriented communication emerge in open networking architectures, such as O-RAN.
Biographies
Shashi Raj Pandey is currently an assistant professor at the Connectivity Section, Aalborg University, Denmark. He received his Ph.D. in computer science and engineering from Kyung Hee University, South Korea, in 2021. His research interests include network economics, wireless communications and networking, distributed machine learning, and semantic communications.
Fabio Saggesse is a postdoctoral researcher in the Connectivity Section at Aalborg University, Denmark. He earned his Ph.D. in Information Engineering from the University of Pisa, Italy, in 2022. In 2024, he received the Marie Skłodowska-Curie Postdoctoral Fellowship. His research focuses on PHY/MAC wireless connectivity, goal-oriented communications, and integrated sensing and communications.
Junya Shiraishi is currently a Marie Skłodowska-Curie Postdoctoral Fellow at the Connectivity Section, Aalborg University, Denmark. He received his Ph.D. in engineering from Kansai University, Japan, in 2023. His research interests include energy-efficient PHY/MAC protocol designs for IoT networks, tiny machine learning for communication, and goal-oriented communication.
Federico Chiariotti is currently an assistant professor at the University of Padova, Italy, where he also received his PhD in information engineering in 2019. He has authored over 100 peer-reviewed papers on latency optimization, semantic and goal-oriented communication, and machine learning in networking, and is currently an Associate Editor of IEEE TWC and IEEE Net. Lett.
Petar Popovski is a Professor at Aalborg University, where he heads the section on Connectivity and a Visiting Excellence Chair at the University of Bremen. He is the Editor-in-Chief of IEEE JSAC. His research interests are in communication theory and wireless connectivity.